OpenStudy (anonymous):

find all numbers for which the rational expression is undefined. s^3-9s/s^2-16

OpenStudy (anonymous):

should just be anything where s^2=0

OpenStudy (anonymous):

with btw is s=0

OpenStudy (anonymous):

which*

OpenStudy (anonymous):

no its undefined when you are dividing something by zero \[\frac{a}{0} \]

OpenStudy (anonymous):

i think the denominator is s^2 - 16, so the answer is +4, -4

OpenStudy (anonymous):

so you want the bottom zero right (s^2 -16) so when does s^2-16 =0 ?